# Deep Learning-Based Beamforming Design Using Target Beam Patterns

### v1

@@ Abstract
Index Terms:
Beamforming, antenna array, beam pattern, deep learning.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) offers transformative advantages in wireless communication by leveraging a large array of antennas at base stations (BSs), dramatically enhancing network capacity and spectral efficiency [1]. Advanced beamforming techniques are widely used in massive MIMO, which allow precise signal directionality, improving signal strength, reducing interference, and extending coverage by concentrating energy in target directions to improve directional gain [2, 3]. However, the difficulty in obtaining accurate channel state information (CSI) due to large-scale antenna arrays and the high operating frequency poses challenges to the system design, as conventional beamforming methods heavily depend on precise CSI to maintain the theoretical performance gains.
@@
Existing beamforming architectures mainly include digital beamforming (DBF), analog beamforming (ABF), and hybrid beamforming (HBF). HBF achieves a balance between hardware cost and spectrum efficiency by jointly optimizing analog and digital domain beamforming matrices. HBF methods include iterative optimization algorithms such as orthogonal matching pursuit (OMP)[4], manifold optimization based alternating minimization (MO-AltMin)[5], as well as methods based on deep learning (DL)[6]. For ABF, beam scanning and selection are usually performed by using predefined codebooks[7]. However, optimal DBF and HBF based on optimization or DL require accurate CSI (i.e., accurate channel matrix) to design the beamforming matrix, causing high overhead in channel estimation. Codebook-based methods require long-term beam training but have limited accuracy due to quantization errors. Array synthesis is a key technology for MIMO beamforming, which utilizes beam patterns to obtain the corresponding beamforming vector. Traditional array synthesis methodologies, including analytical techniques and numerical optimization approaches[8, 9], are effective in small-scale arrays, but their computational complexity rises sharply as the number of array elements increases, making them unsuitable for large-scale MIMO arrays. Existing DL-based array synthesis methods[10, 11] avoid the high computational complexity, but the models are highly dependent on specific array configurations and lack generalization capabilities, changing antennas or scenarios requires re-training of the network. Meanwhile, most of the research focuses on DBF and ABF, but few studies on array synthesis under HBF.
@@
In order to reduce the CSI acquisition overhead in beamforming, we consider using partial CSI for beamforming across diverse antenna array architectural configurations. Specifically, during beamforming, the channel vector between the BS and user equipment (UE) is not available. Instead, only partial information such as the UE’s angular location and beam shape characteristics (e.g., the main lobe direction, shape, and side lobe level) is known. However, such information cannot be directly utilized for the beamforming design based on spectral efficiency maximization. Here we consider using normalized amplitude beam pattern of the channel as the partial CSI, which can be obtained through low-overhead techniques to estimate angles-of-departure (AoDs) and path gain amplitudes[12], fast ray-tracing on 3D maps, or multi-modal-sensing based user localization. A more detailed study of beam pattern acquisition will be discussed in future works. Moreover, customized beam patterns are demanded in scenarios such as directional broadcasting and flat-top coverage. This motivates beamforming approaches that support flexible pattern control while reducing CSI estimation overhead. In this paper, we propose a DL-based beamforming method for all three beamforming architectures using a target beam pattern, which consists of an encoder and a decoder module. Specifically, the encoder extracts low-dimensional features from the complex beam pattern, and then the decoder uses these features to perform beamforming design under each architecture. A two-stage training process is proposed to effectively train the deep neural network (DNN), with pre-training of the encoder and online training for the decoder. Simulation results show the effectiveness of the proposed method in large antenna array synthesis.
@@ II-A System Model
The URA is located in the $yoz$ plane as shown in Fig. 1. Given a beamforming vector $\mathbf{f}$, the two-dimensional far-field beam pattern $\mathbf{X}\in\mathbb{R}^{H\times W}$ is computed via the array factor  [^1] Potential mutual coupling among antennas[14] is not considered herein and is deferred to future work.
@@ III Proposed Method: Deep Learning-Based Beamforming Design With Target Beam Pattern
We propose a DNN for beamforming design based on target beam patterns under arbitrary beamforming architectures. The proposed DNN takes the target beam pattern as input and outputs the corresponding beamforming vector. In this section, we present the architecture, training process, and the design of loss function of the proposed DNN.
@@ III-A Architecture of the Deep Neural Network
The main challenge in obtaining optimal beamforming vectors from target beam patterns lies in extracting vectors from multidimensional matrices, which can be viewed as a feature extraction problem. While convolutional neural networks (CNNs) are suitable for such tasks, practical wireless communication systems face critical constraints including high power consumption, computational demands, and limited storage capacity. Conventional approaches using CNNs with multiple fully connected layers may suffer from excessive parameterization and training difficulties due to the complexity of various beam patterns and the constraints imposed by different beamforming architectures. To address these challenges, a lightweight encoder-decoder structured DNN is proposed as illustrated in Fig. 2. Compared to vision transformer and deep CNNs, which are computationally intensive and data-hungry, the encoder-decoder offers a compact and efficient solution, better capturing local structure in beam patterns with reduced complexity.
The proposed DNN consists of an encoder module and a decoder module. The encoder module extracts essential feature vectors from complex beam patterns; subsequently, the decoder reconstructs beamforming vectors from these encoded features, effectively reducing network parameters and computational complexity. The overall process can be formulated as
@@
For the encoder design, we choose MobileNetV2[15] as the backbone to balance feature extraction capability with computational efficiency. The lightweight architecture encompasses two key innovations: depth-wise separable convolutions and inverted residual blocks. Depth-wise separable convolutions decouple spatial and channel-wise feature processing by first applying single-channel spatial convolutions followed by $1\times 1$ point-wise convolutions, significantly reducing computational complexity. Inverted residual blocks enhance feature representation through an expansion-compression strategy: channel dimensions are first expanded via $1\times 1$ convolutions, followed by depth-wise convolutions for spatial feature extraction, and finally compressed back to lower dimensions. This approach captures complex patterns while maintaining parameter efficiency.
The encoder implementation details are as follows: The input beam pattern undergoes initial convolution through a $3\times 3$ convolution layer with batch normalization and ReLU6 activation. Seven inverted residual blocks are then stacked, with configurations progressively increasing channel dimensions and downsampling spatial resolutions. For instance, the first block uses a 6×6$\times$6 × expansion ratio, 16 output channels, and stride 1. All convolutional layers utilize ReLU6 activation to constrain output ranges for low-precision compatibility. Finally, the regression head, which consists of a $1\times 1$ convolution, global average pooling, and a fully connected layer, is used to produce a fixed-length feature vector.
@@ III-B2 Online training of the decoder
where $\boldsymbol{\Theta}_{e}^{*}$ represents the fixed parameters of the pre-trained $\varepsilon(\cdot)$. As discussed in Section II, the design of an appropriate loss function $\mathcal{L}_{2}$ is critical for beamforming optimization. To meet practical requirements, $\mathbf{X}_{f}$ should exhibit three key characteristics: close alignment with $\mathbf{X}_{T}$ in the main lobe region  [^2] The main lobe region is defined as areas with normalized amplitude exceeding -10 dB in the dB-scale beam pattern., suppressed side lobe  [^3] The side lobe region corresponds to areas with amplitude below -20 dB. levels below those of $\mathbf{X}_{T}$, and similarity in the moderate region (between -20 dB and -10 dB). Accordingly, the composite loss function is formulated as

### v2

@@ I Introduction
Massive multiple-input multiple-output (MIMO) offers transformative advantages in wireless communication by leveraging a large array of antennas at base stations (BSs), dramatically enhancing spectral efficiency [1]. Advanced beamforming techniques are widely used in massive MIMO, which allow for precise signal directionality, improving signal strength, reducing interference, and extending coverage[2, 3]. However, the difficulty in obtaining accurate channel state information (CSI) due to large-scale antenna arrays and the high operating frequency poses challenges to the system design, as conventional beamforming methods heavily depend on precise CSI to maintain the theoretical performance gains.
@@
Existing beamforming architectures mainly include digital beamforming (DBF), analog beamforming (ABF), and hybrid beamforming (HBF). HBF achieves a balance between hardware cost and spectrum efficiency by jointly optimizing analog and digital domain beamforming matrices. HBF methods include iterative optimization algorithms such as orthogonal matching pursuit (OMP)[4], manifold optimization based alternating minimization (MO-AltMin)[5], as well as methods based on deep learning (DL)[6]. For ABF, beam scanning and selection are usually performed by using predefined codebooks[7]. However, optimal DBF and HBF based on optimization or DL require accurate CSI (i.e., accurate channel matrix) to design the beamforming matrix, causing high overhead in channel estimation. Codebook-based methods suffer from quantization errors and require lengthy beam training, while array synthesis is crucial for MIMO beamforming, using beam patterns to derive beamforming vectors. Traditional array synthesis methodologies, including analytical techniques and numerical optimization approaches[8, 9], are effective in small-scale arrays, but their computational complexity rises sharply as the number of array elements increases, making them unsuitable for large-scale MIMO arrays. Existing DL-based array synthesis methods [10, 11] reduce complexity but lack generalization, requiring retraining for new antenna configurations or scenarios. Meanwhile, most of the research focuses on DBF and ABF, but few studies on array synthesis under HBF.
@@
In order to reduce the CSI acquisition overhead in beamforming, we consider using partial CSI for beamforming across diverse antenna array architectural configurations. Specifically, during beamforming, the channel vector between the BS and user equipment (UE) is not available. Instead, only partial information such as the UE’s angular location and beam shape characteristics (e.g., the main lobe direction, shape, and side lobe level) is known. However, such information cannot be directly utilized for the beamforming design based on spectral efficiency maximization. Here we consider using normalized amplitude beam pattern of the channel as the partial CSI, which can be obtained through low-overhead techniques to estimate angles-of-departure (AoDs) and path gain amplitudes[12], fast ray-tracing on 3D maps[13], or multi-modal-sensing based user localization. A more detailed study of beam pattern acquisition will be discussed in future work. Moreover, customized beam patterns are demanded in scenarios such as directional broadcasting[14] and flat-top coverage. This motivates beamforming approaches that support flexible pattern control while reducing CSI estimation overhead. In this paper, we propose a DL-based beamforming method for all three beamforming architectures using a target beam pattern, which consists of an encoder and a decoder module. Specifically, the encoder extracts low-dimensional features from the complex beam pattern, and then the decoder uses these features to perform beamforming design under each architecture. A two-stage training process is proposed to effectively train the deep neural network (DNN), with pre-training of the encoder and online training for the decoder. Simulation results show the effectiveness of the proposed method in large antenna array synthesis.
@@ II-A System Model
The URA is located in the yozyoz plane as shown in Fig. 1. Given a beamforming vector $\mathbf{f}$, the two-dimensional far-field beam pattern $\mathbf{X}\in\mathbb{R}^{H\times W}$ is computed via the array factor  [^1] Potential mutual coupling among antennas[16] is not considered herein and is deferred to future work.
@@ III Proposed Method: Deep Learning-Based Beamforming Design With Target Beam Pattern
We propose a DNN for beamforming design based on target beam patterns. The proposed DNN takes the target beam pattern as input and outputs the corresponding beamforming vector. In this section, we present the architecture, training process, and the design of the loss function of the DNN.
@@ III-A Architecture of the Deep Neural Network
The main challenge in obtaining optimal beamforming vectors from target beam patterns lies in extracting vectors from multidimensional matrices, which can be viewed as a feature extraction problem. While convolutional neural networks (CNNs) are suitable for such tasks, practical wireless communication systems face critical constraints including high power consumption, computational demands, and limited storage capacity. Conventional approaches using CNNs with multiple fully connected layers may suffer from excessive parameterization and training difficulties due to the complexity of various beam patterns and the constraints imposed by different beamforming architectures. To address these challenges, a lightweight encoder-decoder structured DNN is proposed as illustrated in Fig. 2. Compared to computationally intensive and data-hungry Vision Transformers and deep CNNs, the encoder-decoder offers a compact and efficient solution, better capturing local structure in beam patterns with reduced complexity.
The proposed DNN consists of an encoder module and a decoder module. The encoder module extracts essential feature vectors from complex beam patterns, subsequently, the decoder reconstructs beamforming vectors from these encoded features. The overall process can be formulated as
@@
For the encoder design, we choose MobileNetV2[17] as the backbone to balance feature extraction capability with computational efficiency. The lightweight architecture encompasses two key innovations: depth-wise separable convolutions and inverted residual blocks. Depth-wise separable convolutions decouple spatial and channel-wise feature processing by first applying single-channel spatial convolutions followed by $1\times 1$ point-wise convolutions, significantly reducing computational complexity. Inverted residual blocks enhance feature representation through an expansion-compression strategy: channel dimensions are first expanded via $1\times 1$ convolutions, followed by depth-wise convolutions for spatial feature extraction, and finally compressed back to lower dimensions.
The encoder implementation details are as follows: The input beam pattern undergoes initial convolution through a $3\times 3$ convolution layer with batch normalization and ReLU6 activation. Seven inverted residual blocks are then stacked, with configurations progressively increasing channel dimensions and downsampling spatial resolutions. For instance, the first block uses a $6\times$ expansion ratio, 16 output channels, and stride 1. All convolutional layers utilize ReLU6 activation to constrain output ranges for low-precision compatibility. Finally, the regression head, which consists of a $1\times 1$ convolution, global average pooling, and a fully connected layer, is used to produce a fixed-length feature vector.
@@ III-B2 Online training of the decoder
where $\boldsymbol{\Theta}_{e}^{*}$ represents the fixed parameters of the pre-trained $\varepsilon(\cdot)$. As discussed in Section II, the design of an appropriate loss function $\mathcal{L}_{2}$ is critical for beamforming optimization. To meet practical requirements, $\mathbf{X}_{f}$ should exhibit three key characteristics: close alignment with $\mathbf{X}_{T}$ in the main lobe region [^2] The main lobe region is defined as the areas with normalized amplitude exceeding -10 dB in the dB-scale beam pattern., suppressed side lobe [^3] The side lobe region corresponds to the areas with amplitude below -20 dB. levels below those of $\mathbf{X}_{T}$, and similarity in the moderate region (between -20 dB and -10 dB). Accordingly, the composite loss function is formulated as